# Retrieval-Augmented Mixture of LoRA Experts for Uploadable Machine Learning

### v1

@@ 2.3. Personalized LoRA serving
Sheng et al. (2023) propose S-LoRA to discuss serving thousands of concurrent LoRA. The framework targets scenarios in which multiple tasks must be handled simultaneously without compromising the efficiency of the base models. Wen and Chaudhuri (2023) propose FLoRA, which enables efficient batching of diverse request types in the LoRA of foundation models. These studies discuss how to deploy or train personalized LoRAs. However, these methods can only utilize a single user-specified LoRA during inference, failing to fully leverage the combination of LoRAs from different tasks. Moreover, the primary focus of these discussions is on computational strategies in GPUs and training strategies, which are orthogonal to the routing strategies with which we are concerned.
3. Preliminaries
@@
3.1. Low-Rank Adaptation
Directly fine-tuning large language models with all parameters is computationally intensive and is not feasible in low-resource scenarios. Based on the idea that only a small number of low-rank parameters need to be fine-tuned for sufficient performance in new domains, Hu et al. (2021) proposed the Low-Rank Adaptation, where the LoRA module can be combined with the pre-trained parameters in parallel for efficient inference.
@@ 3.1. Low-Rank Adaptation
3.2. Mixture of LoRA Experts
The mixture of LoRA Experts (MoLE) is an efficient ensemble approach that enhances model scalability by selectively activating only a subset of parameters, thus increasing the model’s capacity without incurring additional computational costs. Recently, several studies have explored the integration of Low-Rank Adaptation (LoRA) with MoE for diverse applications. Unlike traditional transformer models, each MoE layer with the original parameter $W$ consists of $k$ independent LoRA experts $\{\Delta W_{i}\}_{i=1}^{k}$, along with a gating function $G(\cdot)$ that models the probability distribution to determine the weighting of these experts’ outputs. The output of a MoLE layer can be given by:
(2) $x^{\prime}=W_{0}x+\sum_{i=1}^{k}G(x)_{i}\cdot\Delta W_{i}x.$
@@ 3.2. Mixture of LoRA Experts
3.3. Problem Formulation of Uploadable Machine Learning
In this part, we give a formal definition of Uploadable Machine Learning. The framework of UML comprises three aspects as shown in Fig.1: upstream Edge-side LoRA contributors, the Multi-LoRA serving system on the service side, and downstream heterogeneous requests.
@@ 3.3. Problem Formulation of Uploadable Machine Learning
The downstream requests are in a mixed-task form, where the mixed task inputs can be formulated as $T_{mix}=\{x\mid x\in T_{1}\lor T_{2}\lor\cdots\lor T_{k}\}$ where $\lor$ stands for the logical disjunction operator. Given the heterogeneous nature of downstream requests, the serving process should provide personalized service for each input prompt.
@@
Figure 2. The RAMoLE Framework. This framework, equipped with a pool of candidate LoRAs from various domains/tasks, is designed to offer personalized services tailored to the input provided. It begins by executing an input-aware LoRA retrieval process aimed at identifying LoRAs corresponding to tasks analogous to the input (§4.1). Subsequently, it employs a specialized LoRA composition mechanism to efficiently utilize the retrieved LoRAs (§4.2). By constructing a LoRA mapping matrix for batch inputs, the framework facilitates effective batch inference (§4.3).
4. Retrieval-Augmented Mixture of LoRA Experts Framework
@@
In this section, we describe the RAMoLE framework as shown in Fig.2 for serving multi-LoRAs in UML. This framework contains three major components: the input-aware LoRA retrieval module (§4.1), the On-the-fly MoLE module (§4.2), and the batch inference strategy (§4.3).
4.1. Input-Aware LoRA Retrieval
@@
To manage a large pool of LoRAs, the initial stage of our framework involves constructing a LoraRetriever specifically designed to efficiently retrieve the appropriate LoRAs for each input, particularly in scenarios where LoRAs are dynamicaly updated. However, existing approaches fall short of accurately identifying LoRAs under such conditions. MoE-based methods  (Wu et al., 2023a; Muqeeth et al., 2023) struggle to generalize when new LoRAs are introduced due to the fixed selection of LoRAs established during router training. Retrieval methods like sentence embedding (Reimers and Gurevych, 2019; Ni et al., 2021) or task embedding (Achille et al., 2019; Zhou et al., 2022) fail to map both samples and LoRA into a shared embedding space, limiting their effectiveness in input-aware LoRA retrieval.
@@
To achieve this goal, we propose to train a retriever via instruction fine-tuning  (Su et al., 2022; Asai et al., 2022), namely LoraRetriever, which can retrieve suitable LoRAs from a massive LoRA pool for a given input sample. The fundamental concept behind LoraRetriever comprises two main steps: (i) First, to embed different task-specific LoRAs into embedding space for facilitating retrieval, we posit that each LoRA can be represented by some data points, which can be obtained by randomly choosing a dozen samples from the training dataset. Then we average their instruction embeddings to represent the embedding of each LoRA. (ii) To improve generalization for unseen LoRAs in LoRA retrieving, we train the retriever through instruction fine-tuning (Su et al., 2022; Wei et al., 2021) on a subset of all tasks. Training on a small subset of tasks is designed to simulate scenarios involving the integration of new LoRAs, thereby underscoring our method’s generalization abilities via instruction fine-tuning. These two strategies enable the effective use of limited data distributions for input-aware retrieval and can be generalized to unseen LoRAs.
@@
Formally, with a sentence-embedding model $E$, input sequence $x$, and the instruction $I$ for embedding purposes, the instructed embedding can be formulated as $E(I\oplus x)$, where $\oplus$ denotes the concatenation operation. In order to allow the embedding to capture the similarity between different tasks, the instruction is expressed as ”Represent the sentence for similar task retrieval”. Each LoRA module is embedded with $m$ randomly selected domain-specific samples, expressed as $E(\phi)=\frac{1}{m}\sum^{m}_{i=1}E(I\oplus x_{i\phi})$. This embedding method integrates both sample-wise and LoRA-module-wise embeddings, facilitating the calculation of similarity between an individual sample and a LoRA module. For measuring the similarity between LoRA module $\phi$ and the input sequence $x$, following  (Ni et al., 2021), we leverage the cosine similarity between the LoraRetriever embeddings: $s(x,\phi,I)=\cos(E(I\oplus x),E(\phi))$.
@@
To improve LoRA retrieval by the retriever and broaden its generalization to unseen LoRAs, we train the embedding model $E$ through instruction fine-tuning on a small subset of tasks. To prevent the need to access new samples, we use previously employed samples for embedding LoRAs as our training data. Consider $t$ distinct training tasks, represented as $\mathcal{T}_{train}=\{T_{1},\cdots,T_{t}\}$. Following Ni et al. (2021), the training dataset $\mathcal{D}$ comprises paired samples ${(x_{i},x_{i}^{+})}$, where each $x_{i}$ is a sample from a task $T_{i}\in\mathcal{T}_{train}$, and a positive sample $x_{i}^{+}$ is randomly selected from the same task $T_{i}$. To complement each positive pair, we randomly select $p$ negative pairs ${(x_{i},x_{ij}^{-})}_{j=1}^{p}$, ensuring that $x_{ij}^{-}$ is sourced from tasks outside of $T_{i}$, thereby $x_{ij}^{-}\notin T_{i}$. The training process is achieved through a contrastive loss (Karpukhin et al., 2020; Izacard et al., 2021; Ni et al., 2021) defined as follows:
@@
4.2. On-the-fly Mixture of LoRA Experts
After retrieving the top-k LoRAs, $\Phi_{i}$, for an input $x_{i}$, we proceed to integrate these LoRAs into the LLM with parameter $\theta$. This section delves into the proposed on-the-fly Mixture of LoRA Experts (DMoLE) mechanism for LoRA composition. We start by introducing the vanilla LoRA composition strategies as described in Zhao et al. (2024), and illustrate the limitations of these vanilla methods. Subsequently, we offer a formal description of the proposed on-the-fly MoLE mechanism, which incorporates an additional LoRA serving as a router, namely RouterLoRA. Its primary purpose is to learn to route among retrieved LoRAs, effectively using cross-attention to assign weights to each one and decouple the routing process from specific LoRA modules.
4.2.1. Vanilla LoRA Composition
@@
Since the LoRAs for different tasks are trained independently, the model is significantly affected by data heterogeneity during the parameter fusion process. This results in catastrophic collapses in performance on respective tasks after fusion due to parameter interference, a phenomenon that becomes more pronounced with the increasing number of composed LoRAs (Zhao et al., 2024). Consequently, the Fusion of LoRAs method is not suitable for application in scenarios involving UML. On the other hand, the mixture of LoRAs method, which averages the outputs of different LoRAs, achieves better performance, especially on unknown tasks. However, because it uniformly assigns the same weight to the output of each LoRA without differentiation, this method fails to fully leverage the most optimal LoRA for the task, leading to a suboptimal model performance.
4.2.2. On-the-fly Mixture of LoRA Experts with RouterLoRA
@@
(9) $s_{i}=\langle q,k_{i}\rangle/\sqrt{r}$
(10) $\alpha=\text{Softmax}(s_{1},s_{2},\ldots,s_{n}).$
@@
(12) $\mathcal{L}_{train}=\frac{1}{|\mathcal{D}_{train}|}\sum_{x,y\in\mathcal{D}_{train}}\mathcal{L}(f(x;\Phi_{train},\gamma,\theta),y).$
@@
(13) $\mathcal{L}_{train}=\frac{1}{|\mathcal{D}_{train}|}\sum_{x,y\in\mathcal{D}_{train}}\mathcal{L}(f(x;Dropout(\Phi_{train},p),\gamma,\theta),y),$
where $p$ denotes the dropout rate. Here, the function $Dropout(\Phi_{train},p)$ selectively deactivates a subset of modules $\Phi_{train}$ at a rate $p$, enabling us to train a more robust model by introducing variability and reducing overfitting.
4.3. Batch Inference of Multiple LoRAs
@@
The LoRA mapping vectors are combined into a matrix $\mathbf{M}\in\mathbb{R}^{b\times p}$. The parameters of a submodule in LoRA can be denoted as $A$ and $B$, and are concatenated within the batched LoRAs $\Phi_{\mathcal{B}}$ to obtain $\mathbf{A}\in\mathbb{R}^{p\times r\times d}$ and $\mathbf{B}\in\mathbb{R}^{p\times d\times r}$. We first gather the LoRA modules for each input within the batch through $\mathbf{A}^{\prime}=\mathbf{M}\otimes\mathbf{A}$ and $\mathbf{B}^{\prime}=\mathbf{M}\otimes\mathbf{B}$ where $\otimes$ denotes the gather operation. By gathering the corresponding parameters for each input, we obtain the matrix $\mathbf{A}^{\prime}\in\mathbb{R}^{b\times k\times r\times d}$ and $\mathbf{B}^{\prime}\in\mathbb{R}^{b\times k\times d\times r}$. Then we can obtain the value matrix $\mathbf{V}$ through $\mathbf{V}=\mathbf{B}^{\prime}\circ\mathbf{A}^{\prime}\circ X$, where $\mathbf{V}\in\mathbb{R}^{k\times b\times l\times d}$ and we extend the symbol $\circ$ to denote potential broadcasting as Wen and Chaudhuri (2023). Then the key matrix is computed through the RouterLoRA in the following way:
(14) $q=A_{r}\circ X$
(15) $\mathbf{K}=B_{R}^{T}\circ\mathbf{V}$
@@
4.4. Evaluation Framework
Figure 3. Datasets and task clusters used to train LoRAs and generate mixed-task evaluation set in this paper (NLU tasks in blue; NLG tasks in green).
5. Experiments
@@
5.0.1. Training And Evaluation Datasets
We leverage a subset of flan-v2 datasets  (Wei et al., 2021) as shown in Fig.3 for LoRA expert training and mixed-task dataset generation. We adopt the task cluster as described by  Wei et al. (2021) and summarize the details of the datasets used as follows:
5.0.2. Base Model & LoRA Configuration.
To test various methods in the mixed-task scenarios, we leverage Llama-2-{7b,13b} (Touvron et al., 2023) as base models and train a range of LoRAs for a spectrum of tasks. We selected a portion of the Flan-v2 datasets (Wei et al., 2021) to train 48 LoRAs for a spectrum of tasks covering Natural Language Understanding (NLU) and Natural Language Generation (NLG). Following the categorization by Wei et al. (2021), these tasks can be grouped into 10 distinct task clusters. We train each LoRA according to the Alpaca (Taori et al., 2023) format and rank $r$, and the scaling hyperparameter $\alpha$ are set to 6 and 12, respectively.

### v2

@@ II-C Personalized LoRA serving
[44] propose S-LoRA to discuss serving thousands of concurrent LoRA. The framework targets scenarios in which multiple tasks must be handled simultaneously without compromising the efficiency of the base models. [45] propose FLoRA, which enables efficient batching of diverse request types in the LoRA of foundation models. These studies discuss how to deploy or train personalized LoRAs. However, these methods can only utilize a single user-specified LoRA during inference, failing to fully leverage the combination of LoRAs from different tasks. Moreover, the primary focus of these discussions is on computational strategies in GPUs and training strategies, which are orthogonal to the routing strategies with which we are concerned.
III Preliminaries
@@
III-A Low-Rank Adaptation
Directly fine-tuning large language models with all parameters is computationally intensive and is not feasible in low-resource scenarios. Based on the idea that only a small number of low-rank parameters need to be fine-tuned for sufficient performance in new domains, [15] proposed the Low-Rank Adaptation, where the LoRA module can be combined with the pre-trained parameters in parallel for efficient inference.
@@ III-A Low-Rank Adaptation
III-B Mixture of LoRA Experts
The mixture of LoRA Experts (MoLE) is an efficient ensemble approach that enhances model scalability by selectively activating only a subset of parameters, thus increasing the model’s capacity without incurring additional computational costs. Recently, several studies have explored the integration of Low-Rank Adaptation (LoRA) with MoE for diverse applications. Unlike traditional transformer models, each MoE layer with the original parameter $W$ consists of $k$ independent LoRA experts $\{\Delta W_{i}\}_{i=1}^{k}$, along with a gating function $G(\cdot)$ that models the probability distribution to determine the weighting of the outputs of these experts. The output of a MoLE layer can be given by:
$x^{\prime}=W_{0}x+\sum_{i=1}^{k}G(x)_{i}\cdot\Delta W_{i}x.$ (2)
@@ III-B Mixture of LoRA Experts
III-C Problem Formulation of Uploadable Machine Learning
In this part, we give a formal definition of Uploadable Machine Learning. The UML framework comprises three aspects as shown in Fig.1: upstream Edge-side LoRA contributors, the Multi-LoRA serving system on the service side, and downstream heterogeneous requests.
@@ III-C Problem Formulation of Uploadable Machine Learning
The downstream requests are in mixed task form, where mixed task input can be formulated as $T_{mix}=\{x\mid x\in T_{1}\lor T_{2}\lor\cdots\lor T_{k}\}$ where $\lor$ represents the logical disjunction operator. Given the heterogeneous nature of downstream requests, the serving process should provide personalized service for each input prompt.
@@
Figure 2: The RAMoLE Framework. This framework, equipped with a pool of candidate LoRAs from various domains/tasks, is designed to offer personalized services tailored to the input provided. It begins by executing an input-aware LoRA retrieval process aimed at identifying LoRAs corresponding to tasks analogous to the input (§IV-A). Subsequently, it employs a specialized LoRA composition mechanism to efficiently utilize the retrieved LoRAs (§IV-B). By constructing a LoRA mapping matrix for batch inputs, the framework facilitates effective batch inference (§IV-C).
IV Retrieval-Augmented Mixture of LoRA Experts Framework
@@
In this section, we describe the RAMoLE framework as shown in Fig.2 for serving multi-LoRAs in UML. This framework contains three major components: the input-aware LoRA retrieval module (§IV-A), the On-the-fly MoLE module (§IV-B), and the batch inference strategy (§IV-C).
IV-A Input-Aware LoRA Retrieval
@@
To manage a large pool of LoRAs, the initial stage of our framework involves constructing a LoraRetriever specifically designed to efficiently retrieve the appropriate LoRAs for each input, particularly in scenarios where LoRAs are dynamicaly updated. However, existing approaches fall short of accurately identifying LoRAs under such conditions. MoE-based methods  [24, 23] struggle to generalize when new LoRAs are introduced due to the fixed selection of LoRAs established during router training. Retrieval methods such as sentence embedding [46, 47] or task embedding [48, 49] fail to map both samples and LoRA into a shared embedding space, limiting their effectiveness in input-aware LoRA retrieval.
@@
To achieve this goal, we propose to train a retriever via instruction fine-tuning  [27, 28], namely LoraRetriever, which can retrieve suitable LoRAs from a massive LoRA pool for a given input sample. The fundamental concept behind LoraRetriever comprises two main steps: (i) First, to embed different task-specific LoRAs into embedding space for facilitating retrieval, we posit that each LoRA can be represented by some data points, which can be obtained by randomly choosing a dozen samples from the training dataset. Then we average their instruction embeddings to represent the embedding of each LoRA. (ii) To improve generalization for unseen LoRAs in LoRA retrieving, we train the retriever through instruction fine-tuning [27, 5] on a subset of all tasks. Training on a small subset of tasks is designed to simulate scenarios involving the integration of new LoRAs, thereby underscoring our method’s generalization abilities via instruction fine-tuning. These two strategies enable the effective use of limited data distributions for input-aware retrieval and can be generalized to unseen LoRAs.
@@
Formally, with a sentence-embedding model $E$, input sequence $x$, and the instruction $I$ for embedding purposes, the instructed embedding can be formulated as $E(I\oplus x)$, where $\oplus$ denotes the concatenation operation. In order to allow the embedding to capture the similarity between different tasks, the instruction is expressed as ”Represent the sentence for similar task retrieval”. Each LoRA module is embedded with $m$ randomly selected domain-specific samples, expressed as $E(\phi)=\frac{1}{m}\sum^{m}_{i=1}E(I\oplus x_{i\phi})$. This embedding method integrates both sample-wise and LoRA-module-wise embeddings, facilitating the calculation of similarity between an individual sample and a LoRA module. For measuring the similarity between LoRA module $\phi$ and the input sequence $x$, following  [47], we leverage the cosine similarity between the LoraRetriever embeddings: $s(x,\phi,I)=\cos(E(I\oplus x),E(\phi))$.
@@
To improve LoRA retrieval by the retriever and broaden its generalization to unseen LoRAs, we train the embedding model $E$ through instruction fine-tuning on a small subset of tasks. To prevent the need to access new samples, we use previously employed samples for embedding LoRAs as our training data. Consider $t$ distinct training tasks, represented as $\mathcal{T}_{train}=\{T_{1},\cdots,T_{t}\}$. Following [47], the training dataset $\mathcal{D}$ comprises paired samples ${(x_{i},x_{i}^{+})}$, where each $x_{i}$ is a sample from a task $T_{i}\in\mathcal{T}_{train}$, and a positive sample $x_{i}^{+}$ is randomly selected from the same task $T_{i}$. To complement each positive pair, we randomly select $p$ negative pairs ${(x_{i},x_{ij}^{-})}_{j=1}^{p}$, ensuring that $x_{ij}^{-}$ is sourced from tasks outside of $T_{i}$, thereby $x_{ij}^{-}\notin T_{i}$. The training process is achieved through a contrastive loss [50, 51, 47] defined as follows:
@@
IV-B On-the-fly Mixture of LoRA Experts
After retrieving the top-k LoRAs, $\Phi_{i}$, for an input $x_{i}$, we proceed to integrate these LoRAs into the LLM with parameter $\theta$. This section delves into the proposed on-the-fly Mixture of LoRA Experts (DMoLE) mechanism for LoRA composition. We start by introducing the vanilla LoRA composition strategies as described in [17], and illustrate the limitations of these vanilla methods. Subsequently, we offer a formal description of the proposed on-the-fly MoLE mechanism, which incorporates an additional LoRA serving as a router, namely RouterLoRA. Its primary purpose is to learn to route among retrieved LoRAs, effectively using cross-attention to assign weights to each one and decouple the routing process from specific LoRA modules.
IV-B1 Vanilla LoRA Composition
@@
Since the LoRAs for different tasks are trained independently, the model is significantly affected by data heterogeneity during the parameter fusion process. This results in catastrophic collapses in performance on respective tasks after fusion due to parameter interference, a phenomenon that becomes more pronounced with the increasing number of composed LoRAs [17]. Consequently, the Fusion of LoRAs method is not suitable for application in scenarios involving UML. On the other hand, the mixture of LoRAs method, which averages the outputs of different LoRAs, achieves better performance, especially on unknown tasks. However, because it uniformly assigns the same weight to the output of each LoRA without differentiation, this method fails to fully leverage the most optimal LoRA for the task, leading to a suboptimal model performance.
IV-B2 On-the-fly Mixture of LoRA Experts with RouterLoRA
@@
$s_{i}=\langle q,k_{i}\rangle/\sqrt{r}$ (9)
$\alpha=\text{Softmax}(s_{1},s_{2},\ldots,s_{n}).$ (10)
@@
$\mathcal{L}_{train}=\frac{1}{|\mathcal{D}_{train}|}\sum_{x,y\in\mathcal{D}_{train}}\mathcal{L}(f(x;\Phi_{train},\gamma,\theta),y).$ (12)
@@
$\mathcal{L}_{train}=\frac{1}{|\mathcal{D}_{train}|}\sum_{x,y\in\mathcal{D}_{train}}\mathcal{L}(f(x;Dropout(\Phi_{train},p),\gamma,\theta),y),$ (13)
where $p$ denotes the dropout rate, here, the function $Dropout(\Phi_{train},p)$ selectively deactivates a subset of modules $\Phi_{train}$ at a rate $p$, enabling us to train a more robust model by introducing variability and reducing overfitting.
IV-C Batch Inference of Multiple LoRAs
@@
The LoRA mapping vectors are combined into a matrix $\mathbf{M}\in\mathbb{R}^{b\times p}$. The parameters of a submodule in LoRA can be denoted as $A$ and $B$, and are concatenated within the batched LoRAs $\Phi_{\mathcal{B}}$ to obtain $\mathbf{A}\in\mathbb{R}^{p\times r\times d}$ and $\mathbf{B}\in\mathbb{R}^{p\times d\times r}$. We first gather the LoRA modules for each input within the batch through $\mathbf{A}^{\prime}=\mathbf{M}\otimes\mathbf{A}$ and $\mathbf{B}^{\prime}=\mathbf{M}\otimes\mathbf{B}$ where $\otimes$ denotes the gather operation. By gathering the corresponding parameters for each input, we obtain the matrix $\mathbf{A}^{\prime}\in\mathbb{R}^{b\times k\times r\times d}$ and $\mathbf{B}^{\prime}\in\mathbb{R}^{b\times k\times d\times r}$. Then we can obtain the value matrix $\mathbf{V}$ through $\mathbf{V}=\mathbf{B}^{\prime}\circ\mathbf{A}^{\prime}\circ X$, where $\mathbf{V}\in\mathbb{R}^{k\times b\times l\times d}$ and we extend the symbol $\circ$ to denote potential broadcasting as [45]. Then the key matrix is computed through the RouterLoRA in the following way:
$q=A_{r}\circ X$ (14)
$\mathbf{K}=B_{R}^{T}\circ\mathbf{V}$ (15)
@@
V Experiments
@@
V-1 Training And Evaluation Datasets
We leverage a subset of flan-v2 datasets  [5] as shown in Fig.9 for LoRA expert training and mixed-task dataset generation. We adopt the task cluster as described by  [5]. The details of datasets can be found in the Appendix. B.
V-2 Base Model & LoRA Configuration.
To test various methods in the mixed-task scenarios, we leverage Llama-2-{7b,13b} [4] as base models and train a range of LoRAs for a spectrum of tasks. We selected a portion of the Flan-v2 datasets [5] to train 48 LoRAs for a spectrum of tasks covering Natural Language Understanding (NLU) and Natural Language Generation (NLG). Following the categorization by [5], these tasks can be grouped into 10 distinct task clusters. We train each LoRA according to the Alpaca [52] format and rank $r$, and the scaling hyperparameter $\alpha$ are set to 6 and 12, respectively.
TABLE I: The average performance of each task cluster. ”IID” means that RAMoLE can access any LoRA for every test sample, encompassing the LoRA specific to the sample’s task. ”OOD” indicates that for each test sample, we mask the LoRA associated with its specific task during the retrieval phase. Consequently, no sample can access its ideal LoRA, allowing us to assess the RAMoLE’s cross-task generalization capability. The performance of perfectly selected corresponding LoRA for each sample is colored in gray. We have bolded the best performance of each task and underlined the best performance in the ”OOD” setting.